At 9.25 percent interest, how long does it take to double your money? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
At 9.25 percent interest, how long does it take to quadruple your money? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

FV = PV(1 + r)t

t = ln(FV / PV) / ln(1 + r)

t = ln(2) / ln(1.0925) = 7.83 years

t = ln(4) / ln(1.0925) = 15.67 years